# Geometric quantum thermodynamics: A fibre bundle approach

### v2

@@ Appendix A Mathematical background
We briefly review the notion of Lie groups and fibre bundles. We focus on the topics that we employed in the main part of the text. The goal is to make the article more self-contained and accessible to researchers interested in quantum thermodynamics, information theory, and gauge theories.
The theory of Lie groups is discussed in Refs. [24, 23], while nice references on fibre bundles are [7, 28, 10]. See also [11, 8] for some topics related to principal bundles. We direct the interested reader to these references for a deeper treatment.
A.1 Bundles and fibre bundles
@@
A bundle is a mathematical structure that generalises the notion of a product space. More specifically, one can understand a bundle as a triple $(\mathds{E},\pi,\mathds{B})$, with $\mathds{E}$ being a topological space called the total space, $\mathds{B}$ a topological space called the base space, and $\pi:\mathds{E}\rightarrow\mathds{B}$ the projection map.
For every point $b\in\mathds{B}$, we call the space $\pi^{-1}(b)$ the fibre of the bundle over $b$. It is often useful to think of a bundle as a set of spaces (fibres) parametrised by the base space and glued together by the topology of the total space.
Bundles are the basic mathematical objects over which many other, more complex structures, such as vector bundles, fibre bundles, and principal bundles, are built. In this work, we are mainly interested in principal bundles. It is, however, often pedagogical to build the notion of a principal bundle from that of a fibre bundle. Thus, we define a fibre bundle by adding two elements to the bundle. A topological space $\mathds{F}$, called the fibre, and a group $\mathrm{G}$ of homeomorphisms of the fibre $\mathds{F}$, called the structure group. Moreover, there is an open covering $\{O_{\alpha}\}$ of $\mathds{B}$ such that, at each $O_{\alpha}$, $\pi^{-1}(O_{\alpha})$ is homeomorphic to the Cartesian product $O_{\alpha}\times\mathds{F}$. The homeomorphisms $\phi_{\alpha}:\pi^{-1}(O_{\alpha})\rightarrow O_{\alpha}\times\mathds{F}$ are called local trivializations of $\mathds{E}$.
Thus, one can see a fibre bundle as a topological space that is locally (but generally not globally) a product space of $\mathds{B}$ and $\mathds{F}$. If a fibre bundle $\xi$ is globally a product space (i.e, if $\mathds{E}=\mathds{B}\times\mathds{F}$), then $\xi=(\mathds{B}\times\mathds{F},\pi,\mathds{F},\mathrm{G},\mathds{B})$ is called a product bundle. A fibre bundle $\zeta=(\mathds{E}^{\prime},\pi^{\prime},\mathds{F},\mathrm{G},\mathds{B})$ is called trivial if it is isomorphic to the product bundle $\xi$.
A section of a fibre bundle is a map $s:\mathds{B}\rightarrow\mathds{E}$ that inverts the projection onto the base space. More specifically, we have $p\circ s=\mathds{1}_{\mathds{B}}$, where $\mathds{1}_{\mathds{B}}$ is the identity map on $\mathds{B}$. The condition that a bundle $\xi$ is trivial is equivalent to the existence of a global section on $\xi$.
@@
Every fibre bundle over a contractible base space $\mathds{B}$ is trivial.
By contractible, we mean that the identity map on $\mathds{B}$ is homotopic to some constant map. In other words, the topological space can be continuously deformed into a single point.
We do not prove this theorem here since it would require many concepts that are outside the scope of this work. For those interested in the proof, Ref. [10] provides an intuitive outline of why Theorem 1 is true, while Ref. [29] provides a rigorous proof using homotopy theory.
A crucial example of a (generally non-trivial) fibre bundle associated with a manifold $\mathds{M}$ is the tangent bundle $\mathrm{T}\mathds{M}$, defined as
@@ A Lie group $\mathrm{G}$ is a group that is also a differentiable manifold in such a way that the map
The Lie algebra $\mathfrak{g}$ associated with a Lie group $\mathrm{G}$ is the space of vector fields invariant under the action of $\mathrm{G}$, with a bracket given by the Lie bracket of vector fields. For our purposes, however, it is more convenient to exploit the fact that the Lie algebra of a Lie group is isomorphic to $\mathrm{T}_{e}\mathrm{G}$, the tangent space of the group at the identity element $e$, and to define the Lie algebra directly as such. Although the Lie algebra is isomorphic to the tangent space at the identity, there is a more general relation between Lie algebras and the tangent spaces of their associated Lie groups, namely:
@@ A.3 Principal Bundles
Principal bundles are among the most important bundles that arise in physics, as they provide the mathematical structure on which one can formalise gauge theories as geometrical constructs.
A principal $\mathrm{G}$-bundle is a fibre bundle for which the structure group (a Lie group $\mathrm{G}$) and the fibre are one and the same. More specifically, given a Lie group $\mathrm{G}$, we define a principal $\mathrm{G}$-bundle $(\mathds{E},\pi,\mathrm{G},\mathds{B})$ as a smooth fibre bundle $(\mathds{E},\pi,\mathds{F},\mathrm{G},\mathds{B})$ such that [23]:
@@
    $\mathrm{G}$ acts freely on the right on $\mathds{E}$;
@@
    The local trivializations $\phi_{\alpha}:\pi^{-1}(O_{\alpha})\rightarrow O_{\alpha}\times\mathrm{G}$ are $\mathrm{G}$-equivariant, where $\mathrm{G}$ acts on $O_{\alpha}\times\mathrm{G}$ by $(b,h)\cdot g=(b,hg)$.
The most important objects in the study of the geometry of a principal $\mathrm{G}$-bundle are connections. More specifically, given a principal $\mathrm{G}$-bundle $\xi=(\mathds{E},\pi,\mathrm{G},\mathds{B})$, we say that a $\mathfrak{g}$-valued $1$-form $\omega$ on $\mathds{E}$ is a connection if and only if
@@
The $A_{p}^{*}$ in the above definition are vector fields on $\mathds{E}$ called the fundamental vector fields. Connections are very important in differential geometry because they allow us to define the notion of parallel transport in general spaces.
Given a connection $\omega$ on a principal $\mathrm{G}$-bundle, one can always construct a curvature form associated with it. That is, from $\omega$, one can make a $\mathfrak{g}$-valued 2-form that acts on vector fields $X$ and $Y$ over the total space $\mathds{E}$ in the following manner
@@
The notion of a connection on a principal $\mathrm{G}$-bundle makes evident the significance of the Maurer-Cartan forms for the geometry of a Lie group. Namely, if we regard a Lie group $\mathrm{G}$ as the principal $\mathrm{G}$-bundle product with total space $\{x\}\times\mathrm{G}$, where $\{x\}$ is a single point, then $\theta_{g}^{R}$ is a connection on the Lie group regarded as a bundle.
@@
Let $\xi=(\mathds{B}\times\mathrm{G},\pi_{1},\mathrm{G},\mathds{B})$ be a trivial principal $\mathrm{G}$-bundle, and $\theta_{g}^{R}$ be the right-invariant Maurer-Cartan form on $\mathrm{G}$. Define the projection $\pi_{2}:\mathds{B}\times\mathrm{G}\rightarrow\mathrm{G}$, $\pi_{2}(x,g)=g$. Then the pullback $\omega_{g}=\pi_{2}^{*}(\theta_{g}^{R})$ is a connection on $\xi$.

### v3

@@ Appendix A Mathematical background
We briefly review the notion of Lie groups and fiber bundles. We focus on the topics we have employed in the main part of the text. The goal is to make the article more self-contained and accessible to researchers interested in quantum thermodynamics, information theory, and gauge theories.
The theory of Lie groups is discussed in Refs. [24, 23], while nice references on fiber bundles are [7, 31, 10]. See also [11, 8] for some topics related to principal bundles. We direct the interested reader to these references for a deeper treatment.
A.1 Bundles and fiber bundles
@@
A bundle is a mathematical structure that generalizes the notion of a product space. More specifically, a bundle can be understood as a triple $(\mathds{E},\pi,\mathds{B})$, with $\mathds{E}$ being a topological space called the total space, $\mathds{B}$ a topological space called the base space, and $\pi:\mathds{E}\rightarrow\mathds{B}$ the projection map.
For every point $b\in\mathds{B}$, we call the space $\pi^{-1}(b)$ the fiber of the bundle over $b$. It is often useful to think of a bundle as a set of spaces (fibers) parametrized by the base space and glued together by the topology of the total space.
Bundles are the basic mathematical objects over which many other, more complex structures, such as vector bundles, fiber bundles, and principal bundles, are built. In this work, we are mainly interested in principal bundles. However, it is often pedagogical to build the notion of a principal bundle from that of a fiber bundle. Thus, we define a fiber bundle by adding two elements to the bundle. A topological space $\mathds{F}$, called the fiber, and a group $\mathrm{G}$ of homeomorphisms of the fiber $\mathds{F}$, called the structure group. Moreover, there is an open cover $\{O_{\alpha}\}$ of $\mathds{B}$ such that, at each $O_{\alpha}$, $\pi^{-1}(O_{\alpha})$ is homeomorphic to the Cartesian product $O_{\alpha}\times\mathds{F}$. The homeomorphisms $\phi_{\alpha}:\pi^{-1}(O_{\alpha})\rightarrow O_{\alpha}\times\mathds{F}$ are called local trivializations of $\mathds{E}$.
Thus, a fiber bundle can be seen as a topological space that is locally (but generally not globally) a product space of $\mathds{B}$ and $\mathds{F}$. If a fiber bundle $\xi$ is globally a product space (i.e, if $\mathds{E}=\mathds{B}\times\mathds{F}$), then $\xi=(\mathds{B}\times\mathds{F},\pi,\mathds{F},\mathrm{G},\mathds{B})$ is called a product bundle. A fiber bundle $\zeta=(\mathds{E}^{\prime},\pi^{\prime},\mathds{F},\mathrm{G},\mathds{B})$ is called trivial if it is isomorphic to the product bundle $\xi$.
A section of a fiber bundle is a map $s:\mathds{B}\rightarrow\mathds{E}$ that inverts the projection onto the base space. More specifically, we have $p\circ s=\mathds{1}_{\mathds{B}}$, where $\mathds{1}_{\mathds{B}}$ is the identity map on $\mathds{B}$. The condition that a bundle $\xi$ is trivial is equivalent to the existence of a global section on $\xi$.
@@
Every fiber bundle on a contractible base space $\mathds{B}$ is trivial.
By contractible, we mean that the identity map in $\mathds{B}$ is homotopic to some constant map. In other words, the topological space can be continuously deformed into a single point.
We do not prove this theorem here, since it would require many concepts that are outside the scope of this work. For those interested in the proof, Ref. [10] provides an intuitive outline of why Theorem 1 is true, while Ref. [32] provides a rigorous proof using homotopy theory.
A crucial example of a (generally non-trivial) fiber bundle associated with a manifold $\mathds{M}$ is the tangent bundle $\mathrm{T}\mathds{M}$, defined as
@@ A Lie group $\mathrm{G}$ is a group that is also a differentiable manifold in such a way that the map
The Lie algebra $\mathfrak{g}$ associated with a Lie group $\mathrm{G}$ is the space of vector fields invariant under the action of $\mathrm{G}$, with a bracket given by the Lie bracket of vector fields. However, for our purposes, it is more convenient to exploit the fact that the Lie algebra of a Lie group is isomorphic to $\mathrm{T}_{e}\mathrm{G}$, the tangent space of the group at the identity element $e$, and to define the Lie algebra directly as such. Although the Lie algebra is isomorphic to the tangent space at the identity, there is a more general relation between Lie algebras and the tangent spaces of their associated Lie groups, namely:
@@ A.3 Principal Bundles
Principal bundles are among the most important bundles that arise in physics, as they provide the mathematical structure on which one can formalize gauge theories as geometrical constructs.
A principal $\mathrm{G}$-bundle is a fiber bundle for which the structure group (a Lie group $\mathrm{G}$) and the fiber are one and the same. More specifically, given a Lie group $\mathrm{G}$, we define a principal $\mathrm{G}$-bundle $(\mathds{E},\pi,\mathrm{G},\mathds{B})$ as a smooth fiber bundle $(\mathds{E},\pi,\mathds{F},\mathrm{G},\mathds{B})$ such that [23]:
@@
    $\mathrm{G}$ acts freely on the right in $\mathds{E}$;
@@
    Local trivializations $\phi_{\alpha}:\pi^{-1}(O_{\alpha})\rightarrow O_{\alpha}\times\mathrm{G}$ are $\mathrm{G}$-equivariant, where $\mathrm{G}$ acts on $O_{\alpha}\times\mathrm{G}$ by $(b,h)\cdot g=(b,hg)$.
The most important objects in the study of the geometry of a principal $\mathrm{G}$-bundle are connections. More specifically, given a principal $\mathrm{G}$-bundle $\xi=(\mathds{E},\pi,\mathrm{G},\mathds{B})$, we say that a $\mathfrak{g}$-valued $1$-form $\omega$ in $\mathds{E}$ is a connection if and only if
@@
The $A_{p}^{*}$ in the above definition are vector fields on $\mathds{E}$ called fundamental vector fields. Connections are very important in differential geometry because they allow us to define the notion of parallel transport in general spaces.
Given a connection $\omega$ on a principal $\mathrm{G}$-bundle, one can always construct a curvature form associated with it. That is, from $\omega$, one can make a $\mathfrak{g}$-valued 2-form that acts on the vector fields $X$ and $Y$ over the total space $\mathds{E}$ in the following manner:
@@
The notion of a connection on a principal $\mathrm{G}$-bundle makes evident the significance of the Maurer-Cartan forms for the geometry of a Lie group. That is, if we consider a Lie group $\mathrm{G}$ as the principal $\mathrm{G}$-bundle product with total space $\{x\}\times\mathrm{G}$, where $\{x\}$ is a single point, then $\theta_{g}^{R}$ is a connection on the Lie group regarded as a bundle.
@@
Let $\xi=(\mathds{B}\times\mathrm{G},\pi_{1},\mathrm{G},\mathds{B})$ be a trivial principal $\mathrm{G}$-bundle, and $\theta_{g}^{R}$ be the right-invariant Maurer-Cartan form on $\mathrm{G}$. Define projection $\pi_{2}:\mathds{B}\times\mathrm{G}\rightarrow\mathrm{G}$, $\pi_{2}(x,g)=g$. Then the pullback $\omega_{g}=\pi_{2}^{*}(\theta_{g}^{R})$ is a connection on $\xi$.